OpenStudy (anonymous):

find the equation of the tangent line to the graph of the given function at the given value x posted below

OpenStudy (anonymous):

\[x^3\sqrt{x^3+8};x=1\]

OpenStudy (anonymous):

find the derivative because derivative= slope then plug in your x value to get m or slope.

OpenStudy (anonymous):

alright thanks man

OpenStudy (anonymous):

3x^2 squared would just be 3x so you get 3x^(2)*3x. You know how to use the power rule right?

OpenStudy (anonymous):

yeah i know power rule, quotient rule, ect...

OpenStudy (anonymous):

why need power rule if you need to find the derivative and plug it in.

OpenStudy (anonymous):

ok cool so using the power rule you get (3x^2)(3x) after plugging in x=1 you just get 9 so slope= 9 now you just need to find the y intercept y=mx+b

OpenStudy (anonymous):

to do that you plug x=1 into original function, in order to find y.

OpenStudy (anonymous):

then finally you plug in your y value, x value, and slope to calculate b or y intercept

OpenStudy (anonymous):

let me know if you get stuck bro @CalcDerp102

OpenStudy (anonymous):

so i know how to do this, but im confused where power rule comes into place i did the derivative and i get 27 + 1/2 but i know its not that

OpenStudy (anonymous):

wait nvm i got it, thanks man! its y = 19/2x-13/2